### v2

# Locally Convex Global Loss Network for Decision-Focused Learning

Abstract
In decision-making problem under uncertainty, predicting unknown parameters is often considered independent of the optimization part. Decision-focused Learning (DFL) is a task-oriented framework to integrate prediction and optimization by adapting predictive model to give better decision for the corresponding task. Here, an inevitable challenge arises when computing gradients of the optimal decision with respect to the parameters. Existing researches cope this issue by smoothly reforming surrogate optimization or construct surrogate loss function that mimic task loss. However, they are applied to restricted optimization domain. In this paper, we propose Locally Convex Global Loss Network (LCGLN), a global surrogate loss model which can be implemented in a general DFL paradigm. LCGLN learns task loss via partial input convex neural network which is guaranteed to be convex for chosen inputs, while keeping the non-convex global structure for the other inputs. This enables LCGLN to admit general DFL through only a single surrogate loss without any sense for choosing appropriate parametric forms. We confirm effectiveness and flexibility of LCGLN by evaluating our proposed model with three stochastic decision-making problems.
1 Introduction
Figure 1: A model training pipeline for PFL, DFL and surrogate DFL. PFL trains predictive model by minimizing the prediction loss. DFL directly delivers gradients minimizing the task loss. Surrogate DFL first learns a surrogate loss model that follows the true task loss well by sampling predictions and its task losses. Then, it trains the predictive model conveying useful gradients in an end-to-end manner.
Decision-making problems under uncertainty arise in various real-world applications such as production optimization, energy planning, and asset liability management (Shiina and Birge 2003; Carino et al. 1994; Fleten and Kristoffersen 2008; Delage and Ye 2010; Garlappi, Uppal, and Wang 2007). These problems involve two main tasks: Prediction and Optimization. The Prediction task aims to create a model for uncertainty and estimate unknown parameters from some input features. This task is often performed using various machine learning (ML) techniques such as regression or deep learning. In the Optimization task, the corresponding optimization problem is solved using diverse off-the-shelf solvers with the estimated parameters from the prediction task. For example, in asset liability management, asset returns are estimated in the prediction stage, and the optimal portfolio to repay the liability is obtained in the optimization stage based on these predictions.
Prediction-focused learning (PFL) framework is the most commonly used approach to solve such problems, handling the two tasks in separate and independent steps. PFL first trains ML model on the input features to produce predictions that closely match the observed parameters. Subsequently, the decision-making problem is defined using the parameters predicted by the trained ML model and is solved to obtain an optimal decision. PFL is based on the underlying belief that accurate predictions will lead to good decisions. However, ML models frequently fail to achieve perfect accuracy, resulting in sub-optimal decisions. Consequently, in many applications, the prediction and optimization tasks are not distinct but are intricately linked and should ideally be considered jointly.
Decision-focused learning (DFL) achieves the above purpose by directly training ML model to make predictions that lead to good decisions. DFL combines prediction and optimization tasks by creating an end-to-end system that aims to directly minimize task loss. Task loss is defined as the quality of decisions derived from the predictive model and therefore, depends on the optimal solution of the associated optimization problem. To train an ML model in this context, differentiating through the optimization problem is required to calculate the gradient of the task loss. This presents a key challenge in integrating prediction and optimization. However, differentiation may be impossible if the decision variable is discrete or if the objective function is discontinuous.
While surrogate optimization still necessitates differentiation of the optimization problem, efforts have been made to develop solver-free surrogate loss functions that can effectively obtain gradients. We refer to this approach as Surrogate DFL. Recent successful works in Surrogate DFL include SPO (Elmachtoub and Grigas 2022), Contrastive Loss (Mulamba et al. 2020), LTR Loss (Mandi et al. 2022), LODL (Shah et al. 2022), EGL (Shah et al. 2024), LANCER (Zharmagambetov et al. 2024), and TaskMet (Bansal et al. 2024). SPO, Contrastive loss, and LTR loss are specifically designed for linear objectives while LODL can be implemented for general DFL problems. However, LODL creates and trains a surrogate with a specific parametric form for each data instance resulting in expensive computation, with decision quality heavily dependent on the chosen parametric surrogate form. While EGL and LANCER have partially addressed the computational cost issue by extending a local model to a global model, the challenge of selecting an appropriate parametric surrogate form still remains. Alongside the surrogates, TaskMet minimized the prediction error while maintaining the decision quality. We illustrated the overall procedures of PFL, DFL, and Surrogate DFL in Figure 1.
Notably, there have been attempts to construct surrogate loss functions using neural networks (Shah et al. 2022). However, these approaches generally performed poorly in experiments because naive MLPs fail to capture the local characteristics of the true underlying loss.
In this paper, we propose Locally Convex Global Loss Network (LCGLN), which is a global and general surrogate loss for DFL. LCGLN adopts partial input convex neural network (PICNN) (Amos, Xu, and Kolter 2017) as a parametric surrogate form to approximate task loss. With PICNN, we guarantee the surrogate loss function to be convex for the chosen inputs, while maintaining the general structure for the others. Furthermore, users are not required to possess an artistic sense for selecting the suitable parametric function form for the task loss. Therefore, LCGLN can handle the general DFL problems using only one surrogate loss, regardless of the number of observed data. We demonstrate the capability of LCGLN on three stochastic decision-making problem, namely inventory stock problem, budget allocation problem and portfolio optimization problem and show that LCGLN learns task loss well with a single surrogate loss.
2 Related Works
Various methodologies based on gradient based learning in DFL have been developed (Mandi et al. 2023). Some approaches directly differentiate the constrained optimization problem to update the model parameters. For instance, stochastic optimization problems were tackled by directly differentiating convex QPs using KKT optimality conditions and employing Optnet (Amos and Kolter 2017) for efficient differentiation (Donti, Amos, and Kolter 2017). For general convex optimization problems, the differentiable solver Cvxpylayers (Agrawal et al. 2019) were developed.
Alternatively, some works introduce regularization terms to smooth the optimization mapping. For example, the Euclidean norm of the decision variables was added to use quadratic programming methods (Wilder, Dilkina, and Tambe 2019), while the logarithmic barrier term was added to differentiate linear programming problems (Mandi and Guns 2010). Furthermore, entropy terms were incorporated to solve multi-label problems (Martins and Kreutzer 2017; Amos, Koltun, and Kolter 2019), while constrained optimization problems were addressed using the perturb-and-MAP (Papandreou and Yuille 2011) framework, which added regularization through perturbations (Niepert, Minervini, and Franceschi 2021; Berthet et al. 2020).
Recent works on DFL have focused on constructing differentiable surrogate loss models. Smart ”Predict, Then Optimize” (SPO) (Elmachtoub and Grigas 2022) proposed a convex SPO+ loss where the loss upper bounds the task loss. To update the predictive model, they obtained a subgradient of their proposed surrogate loss. NCE (Mulamba et al. 2020) used a noise contrastive approach (Gutmann and Hyvärinen 2010) to build a family of surrogate loss for linear objectives. LTR (Mandi et al. 2022) applied this approach on ranking problems. SO-EBM (Kong et al. 2022) used an energy-based differentiable layer to model the conditional probability of the decision. The energy-based layer plays as a surrogate of the optimization problem. LODL (Shah et al. 2022) used supervised machine learning to locally construct surrogate loss models to represent task loss. They first sampled predictions near true instances for each instance and built a surrogate loss model respectively. They proposed three families for the model (e.g. WeightedMSE, Quadratic) to design a parametric surrogate loss models. EGL (Shah et al. 2024) extends LODL to global surrogate model and LANCER (Zharmagambetov et al. 2024) learns such handcraft global surrogate loss via actor-critic framework.
3 Preliminaries
In this section, we motivate necessity of DFL approach from simple knapsack example. We also briefly summarize surrogate loss models and input convex neural networks.
Figure 2: A simple example of knapsack problem. There are two items valued \$40, \$30 each, marked with yellow star. We predict the value of items and choose the higher one. Blue dots and red crosses are predicted values representing good and bad decisions respectively. PFL gives same prediction loss for every prediction while DFL gives \$10 loss in red cross and 0 in blue.
Comparison on PFL and DFL
Prediction focused learning (PFL) and decision focused learning (DFL) are two big learning pipelines for the decision making problems under uncertainty as in Figure 1. Decision making problems can be considered in two big stages: model training stage and evaluation stage. We would like to focus on the model training stage. PFL learns the predictive model focusing on the prediction from the predictive model while DFL focuses on the objective (or task loss for minimization). Starting with PFL, the model training stage is again divided in two steps, prediction and optimization. In prediction step, PFL learns a predictive model by minimizing the prediction loss such as MSE. In optimization step, it solves a constrained optimization problem using the prediction of predictive model. By contrast, the predictive model in DFL is trained in an end-to-end manner. It focuses on learning the model to make good decisions (or actions), and therefore it uses task loss (or objective for maximization problem), rather than concentrating on minimizing the prediction loss.
To give a motivation in usage of DFL, we introduce a motivating example with a simple knapsack problem in Figure 2. Let’s consider two items: item 1 with value \$40 and item 2 with \$30 (true value, yellow star). Assume we can only select one item. Our objective is to predict the value of items by only observing the item features and choose one with a higher value. If we decide to choose item 1 (the optimal decision in this case), we would be happier, earning \$40. Choosing item 2 is a sub-optimal decision, leading to a relative \$10 loss. Suppose we predicted the values as blue dots and red crosses. In PFL point of view, the prediction loss of our prediction loss is equal to \$20. Any prediction on the grey circle will give same prediction loss, which is not informative. In DFL standpoint, blue dots return zero task loss while red crosses gives task loss of \$10. We show that minimizing the prediction loss may not lead to minimizing the task loss.
Surrogate DFL
DFL compute task loss gradients with respect to the input parameters by for example, directly differentiating the optimization problem (or optimizer) (Donti, Amos, and Kolter 2017). However, the optimizer cannot be differentiated easily in most of the general cases. To tackle this, methods using surrogate loss functions were proposed (Elmachtoub and Grigas 2022). Surrogate loss models are differentiable and it approximates the mapping between the prediction and the task loss. As a consequence, it can be used to calculate the gradient to update the predictive model efficiently.
Given dataset $\mathcal{D}=\{(x_{i},y_{i})\}^{N}_{i=1}$ with $N$ instances, our goal is to minimize the regret $\mathcal{R}$ for the optimization problem:
$\min_{\theta}$ $\quad\mathcal{R}(\hat{y},y)$ :$\vcentcolon=\mathcal{L}_{task}(a^{*}(\hat{y}),y)-\mathcal{L}_{task}(a^{*}(y),y)$ (1)
s.t. $\quad\hat{y}$ $=\mathcal{M}_{\theta}(x)$
$a^{*}(\hat{y})$ $=\operatorname*{arg\,min}_{a\in\mathcal{A}}\mathcal{L}_{task}(a,\hat{y})$
where $\mathcal{L}_{task}$ is the task loss to be optimized, $\hat{y}$ is the prediction from the predictive model $\mathcal{M}$ parameterized by $\theta$ and $a^{*}\in\mathcal{A}$ is an optimal action (or decision) in a feasible region $\mathcal{A}$ derived by any off-the-shelf solver. Note that the second term of the objective is nothing but a constant optimal loss, and therefore our objective is equivalent to minimizing $\mathcal{L}_{task}(a^{*}(\hat{y}),y)$.
To update the predictive model via gradient descent w.r.t. $\theta$, one must find the gradient $\partial\mathcal{R}/\partial\theta$. Using the chain rule, the gradient can be decomposed to:
$\frac{\partial\mathcal{R}(\hat{y},y)}{\partial\theta}$ $=\frac{\partial\mathcal{R}(\hat{y},y)}{\partial\hat{y}}\cdot\frac{\partial\hat{y}}{\partial\theta}$ (2)
$\approx\frac{\partial\mathcal{L}_{\psi}(\hat{y},y)}{\partial\hat{y}}\cdot\frac{\partial\hat{y}}{\partial\theta}$ (3)
where $\mathcal{L}_{\psi}(\hat{y},y)$ is the parametric surrogate loss with parameter $\psi$ that is trained to approximate $\mathcal{R}(\hat{y},y)$. The global surrogate loss can be trained to richly approximate the true task loss as:
$\psi^{*}=\operatorname*{arg\,min}_{\psi}\mathop{\mathbb{E}}_{\hat{y},y}[\lVert\mathcal{L}_{\psi}(\hat{y},y)-\mathcal{R}(\hat{y},y)\rVert]$ (4)
Training the predictive model $\mathcal{M}_{\theta}$ can be easily done with backpropagation using the gradient $\partial\mathcal{L}_{\psi}/\partial\theta$.
Partial Input Convex Neural Network
Amos et al. (Amos, Xu, and Kolter 2017) proposed input convex neural network (PICNN) which ensures convexity with respect to the chosen inputs. PICNN is constructed by introducing additional weights to connect input layer to each hidden layer. Given this structure, a non-decreasing convex activation functions such as softplus and non-negativity of weights connecting between hidden layers are required to guarantee convexity of PICNN. LCGLN uses PICNN to richly represent the non-convex nature of the true loss mapping, and simultaneously ensure local convexity.
4 Locally Convex Global Loss Network
In various optimization problems, each has its own specific objectives known as task losses. Our goal is to devise a surrogate loss function capable of accurately representing the true task loss across a range of optimization problems. In this paper, we introduce LCGLN, where a single loss representation can replace the true task loss. While DFL methods suffer from differentiating through the optimization solver when the optimization problem is not smooth, our LCGLN is easily differentiable, and therefore can be easily back-propagated when updating the predictive model via gradients.
4.1 End-to-end Training Procedure using LCGLN
Algorithm 1 Training Predictive Model with LCGLN $\mathcal{L}_{\psi}$
  Dataset: $\mathcal{D}=\{(x_{i},y_{i})\}^{N}_{i=1}$
  1. Generate Samples:
  Sample set $\mathcal{S}\leftarrow\emptyset$
  for $i=1,\ldots,N$ do
     $\mathcal{S}\leftarrow\mathcal{S}\cup(y_{i},y_{i},0)$
     Initialize sampling model $\mathcal{M}_{\xi}$.
     for  $k=1,\ldots,K-1$  do
        $\tilde{y}_{i}^{(k)}=\mathcal{M}_{\xi}(x_{i})$
        $\xi\leftarrow$ Update$(\xi,\nabla_{\xi}\lVert\tilde{y}_{i}^{(k)}-y_{i}\rVert)$
        Solve $a^{*(k)}_{i}=\operatorname*{argmin}_{a}\mathcal{L}_{task}(a(\tilde{y}_{i}^{(k)}),y_{i})$.
        $\mathcal{S}\leftarrow\mathcal{S}\cup(\tilde{y}_{i}^{(k)},y_{i},\mathcal{R}(\tilde{y}_{i}^{(k)},y_{i}))$
     end for
  end for
  2. Learn LCGLN $\mathcal{L}_{\psi}$:
  Initialize surrogate loss model $\mathcal{L}_{\psi}$.
  for $(\hat{y},y,\mathcal{R}(\hat{y},y))$ in $\mathcal{S}$ do
     $\psi\leftarrow$ Update $(\psi,\nabla_{\psi}\lVert\mathcal{L}_{\psi}(\hat{y},y)-\mathcal{R}(\hat{y},y)\rVert)$
  end for
  3. Train Predictive Model $\mathcal{M}_{\theta}$:
  Initialize predictive model $M_{\theta}$.
  $\theta\leftarrow$ Update $(\theta,\nabla_{\theta}\hat{y}\cdot\nabla_{\hat{y}}\mathcal{L}_{\psi}(\hat{y},y))$
Here, we introduce the end-to-end training procedure for LCGLN. The goal is to learn a regret mapping $(\hat{y},y)\rightarrow\mathcal{R}(\hat{y},y)$ to obtain gradient $\partial\mathcal{R}/\partial\theta$ for updating the predictive model $\mathcal{M}_{\theta}$.
Generating Samples
To train the LCGLN in a supervised learning manner, we generate $K$ samples for each $N$ instances, i.e. for each instance $y_{i}$, we sample $\tilde{y}_{i}^{(1)},\tilde{y}_{i}^{(2)},...,\tilde{y}_{i}^{(K)}$.
Previous research has assumed that predictions $\tilde{y}$ will closely approximate the true labels $y$, typically generating sample predictions by simply adding Gaussian noise. To capture realistic predictions, we adopt the model-based sampling (MBS) approach (Shah et al. 2024). This involves constructing a sampling model $M_{\xi}$ that mirrors the architecture of the predictive model and training within a PFL paradigm using MSE. During the training, we take inferences to the intermediate model and generate samples.
Learning Global Surrogate Loss LCGLN
Our objective is to learn a mapping $(\hat{y},y)\rightarrow\mathcal{R}$. For the loss model to convey informative gradients, we claim that $\mathcal{L}_{\psi}$ should be:
  • Expressive: Good surrogate loss model should have sufficient number of parameters to accurately approximate the true task loss. A lack of expressiveness may highly lead to under-performance. We leverage neural networks to ensure sufficient expressiveness.
  • Easily Differentiable: We need to differentiate our loss model with respect to $\hat{y}$ for gradients training the predictive model. Discontinuous or non-differentiable loss models will highly likely to suffer in such task. We wish our model to be easily differentiable.
  • Locally Convex, but not Globally: We want the model to capture the highly non-convex structure of the true underlying loss mapping. At the same time, we claim the loss model should be locally convex for every instance $y_{i}$, where the local minima is formed when $\hat{y}=y$. We induce the local minima by adding $\{(y_{i},y_{i},0)\}_{i=1}^{N}$ to the dataset for every instance.
Table 1: The table contains normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ with standard error mean (SEM) tested on three stochastic optimization problems. The metric is lower the better with 0 meaning optimal. We bold-lettered the best performing results for each problem. We evaluate PFL, DFL, local and global surrogate loss models. We use 32 samples for the surrogate loss models. The global surrogate loss LCGLN outperformed the baselines across all three problems.
Paradigm Methods Problem
Inventory Budget (500 Fakes) Portfolio
2 Stage PFL 0.242 $\pm$ 0.005 0.513 $\pm$ 0.016 0.189 $\pm$ 0.002
Exact Diff DFL 0.228 $\pm$ 0.002 0.532 $\pm$ 0.020 0.187 $\pm$ 0.002
Local Surrogate LODL-DQ 0.378 $\pm$ 0.007 0.503 $\pm$ 0.020 0.193 $\pm$ 0.002
Global Surrogate LANCER 0.182 $\pm$ 0.004 0.490 $\pm$ 0.010 0.246 $\pm$ 0.008
EGL-WMSE 0.371 $\pm$ 0.002 0.510 $\pm$ 0.013 0.187 $\pm$ 0.001
EGL-DQ 0.369 $\pm$ 0.007 0.492 $\pm$ 0.005 0.256 $\pm$ 0.002
LCGLN 0.174 $\pm$ 0.002 0.468 $\pm$ 0.009 0.185 $\pm$ 0.000
Gathering above approaches, we finally propose our Locally Convex Global Loss Network (LCGLN), a neural network that optimizes over the convex inputs to the network and outputs the regret. LCGLN uses partial input convex neural network (Amos, Xu, and Kolter 2017) to learn surrogate loss convex to the designated inputs with sufficient number of parameters. For LCGLN, we do not need to choose the parametric loss form to be used; we can simply use a single LCGLN to replace task loss. Furthermore, we can easily differentiate the model by using a built-in backward functions and obtaining gradients. Most importantly, by constructing a surrogate loss convex in each instance $y_{i}$ while capturing the non-convex nature among instances, LCGLN can richly represent the true task loss.
Training Predictive Model $M_{\theta}$
The predictive model learns a mapping $x\rightarrow y$. We train our predictive model $M_{\theta}$ via gradient descent using the gradients provided by the global surrogate loss $\mathcal{L}_{\psi}$. The procedure from generating dataset to training the predictive model is summarized in Algorithm 1.
5 Experiments and Results
We validate our methodology with three stochastic optimization problems.
5.1 Experimental Settings
Here we explain the experimental details and the baselines used to evaluate our method. Each problem is explained in two stages: the parameter prediction stage and the optimization stage.
Figure 3: A line plot showing the normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ for each methodology and problem setting. The metric is lower the better, with 0 representing the optimal value. We tested sample size if $\{2,4,8,16,32\}$ for each problem. The standard error mean for each experiment is detailed in Appendix C.2. Our global surrogate loss LCGLN represented by the red straight line outperforms when 32 samples are used.
Problem Description
We conducted experiments on three different stochastic optimization problems, each presenting unique challenges as explored in previous researches (Donti, Amos, and Kolter 2017; Wilder, Dilkina, and Tambe 2019; Shah et al. 2022). Our experiments were built on top of the public codes from previous researches (Donti, Amos, and Kolter 2017; Shah et al. 2022). Further details on problem formulation can be found in Appendix B.
  • Inventory Stock: The objective of inventory stock problem (Donti, Amos, and Kolter 2017) is to decide the order quantity $a$ to minimize the cost over the stochastic demand $y$. We generate problem instances from randomly sampled $x\in\mathbb{R}^{20}$ with the non-linear mapping $p(y|x;\theta)\propto\text{exp}(\Theta^{\top}x)$ for the fixed true parameter $\Theta\in\mathbb{R}^{20\times 5}$. We assume the demands are discrete with values $d_{1},d_{2},...,d_{5}$ as $1,2,5,10,20$ with probabilities $(p_{\theta})_{i}\equiv p(y=d_{i}|x;\theta)$.
    Note that the realization $y$ of $p_{\theta}$ is known while the true distribution $p_{\theta}$ is unknown; similar to the real-world settings. Since, the true distribution $p_{\theta}$ is unknown, we transform $y$ to a $p_{\theta}$ form when training the surrogate loss functions.
    Prediction: We predict $p_{\theta}$, the distribution of the demand $y$.
    Optimization: Given the demand distribution $p_{\theta}$, we minimize the cost:
    $\min_{a\in\mathbb{R},\quad a_{b},a_{h}\in\mathbb{R}^{k}}$ $c_{0}a+\frac{1}{2}q_{0}a^{2}+\sum_{i=1}^{k}(p_{\theta})_{i}\biggl{(}c_{b}(a_{b})_{i}$
    $+\frac{1}{2}q_{b}(a_{b})_{i}^{2}+c_{h}(a_{h})_{i}+\frac{1}{2}q_{h}(a_{h})_{i}^{2}\biggr{)}$
    s.t. $d-a\boldsymbol{1}\leq a_{b},$
    $a\boldsymbol{1}-d\leq a_{h},$
    $a,a_{h},a_{b}\geq 0.$
  • Budget Allocation: The objective of the submodular optimization problem (Wilder, Dilkina, and Tambe 2019) is to choose $B=2$ websites to advertise based on click-through rates (CTRs) of $U=10$ users among $W=5$ websites. We generate the user features $x=Ay$ multiplying a random matrix $A$ to the CTRs $y$. We conduct experiments with $\{0,5,50,500\}$ fake targets to add difficulty by concatenating fake target size of random CTRs to the original CTRs.
    Prediction: Predict the CTRs $\hat{y}_{w}$ for $W$ websites and $U$ users where user features $x_{u}$ for website $u$ are given.
    Optimization: Choose $B$ websites to advertise that maximizes the expected number of users that click on the advertisement at least once:
    $\max_{a\in\mathbb{R}^{W}}$ $\sum_{u=1}^{U}(1-\prod_{w=1}^{W}a_{w}\cdot\hat{y}_{wu})$
    s.t. $a\in\{0,1\}^{W}$
  • Portfolio Optimization: We choose the portfolio optimization problem to test when good prediction leads to good decisions. We use historical daily stock return data from 2004 to 2017 for 50 stocks (Shah et al. 2022).
    Prediction: Given historical daily stock return data, predict the future stock price $y$.
    Optimization: Given correlation matrix $\Sigma$, decide a portfolio weight $a$ where we want to maximize the expected return for a given level of risk:
    $\max_{a\in\mathbb{R}^{50}}$ $y^{\top}a-\gamma\cdot a^{\top}\Sigma a$
    s.t. $\sum_{i=1}^{50}a_{i}=1$
    where the risk aversion constant $\gamma=0.1$.
Figure 4: A histogram presenting normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ with standard error mean (SEM) for global surrogate loss models in budget allocation with varying number of fake targets. We use 16 samples for learning loss. The metric is lower the better and 0 when optimal. We test with $\{0,5,50,500\}$ fake targets, noting that the problem becomes more challenging as the number of fake targets increases. Our LCGLN shown in red bars outperforms for most settings.
Baselines
We compare our global surrogate loss network LCGLN with the following baselines from previous researches. When available, we utilized the hyperparameters specified in previous research for the baselines. In cases where these were not provided, we endeavored to fine-tune to perform their best.
  • PFL: A standard approach that trains predictive model on the input features to produce predictions that closely match the observed parameters. We use the negative log-likelihood (NLL) loss for the inventory stock problem and the mean squared error (MSE) loss for the other problems.
  • DFL: A direct approach that differentiate through the solver to derive gradients for training the predictive model. For inventory stock and portfolio optimization problem, we use differentiable QP solver (Donti, Amos, and Kolter 2017). We use multilinear relaxation (Wilder, Dilkina, and Tambe 2019) for the budget allocation problem.
  • LODL: The local surrogate model (Shah et al. 2022) that learns a surrogate loss locally for each instance, creating total $N$ surrogate loss models. To learn a loss model, LODL generates $K$ number of samples for each $N$ instances and solves for the input of the model. We use directed-quadratic surrogate loss (LODL-DQ) from their paper, where it shows the most promising result among the proposed 4 different loss models.
  • LANCER: The global surrogate model (Zharmagambetov et al. 2024) that uses alternating optimization to learn the predictive model and the global surrogate loss model. The prediction samples are sampled from the predictive model and is used to train the surrogate loss. The predictive model is then trained using the gradients from the updated surrogate loss.
  • EGL: The global surrogate model (Shah et al. 2024) that learns a global surrogate loss with feature-based parameterization and model-based sampling. We use directed-quadratic (EGL-DQ) and weighted-MSE (EGL-WMSE) for the baseline, which showed the best results on budget allocation and portfolio optimization problem respectively.
Evaluation Metric
We use the normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ as the evaluation metric, where $\mathcal{R}_{test}$ is a test regret with regret defined in Equation 1. For the budget allocation and portfolio optimization problem which are the maximization task, the regrets are calculated by multiplying the negative sign on the corresponding objective. We calculate the worst case regret $\mathcal{R}_{worst}$ for each problem and derive the normalized test regret. For the inventory stock problem, we assumed the worst-case scenario when the company ordered no stock. In the budget allocation problem, the worst-case scenario was assumed when the advertisements were allocated to the lowest predicted CTRs. For the portfolio optimization problem, we considered the worst-case scenario to be when the entire investment was allocated to the stock with the lowest predicted return. Our metric is 0 when optimal and 1 when worst.
We use a predictive model $\mathcal{M}_{\theta}$ as one hidden layer MLP and a learning rate of 0.001. 500 hidden nodes were used for the portfolio optimization and 10 for the other problems. We mirrored the predictive model exactly for the sampling model $\mathcal{M}_{\xi}$. For LCGLN, we employ a single hidden layer ICNN with two nodes per layer, a learning rate of 0.001 and a softplus activation function. A figure illustrating the LCGLN can be found in Appendix A.1. For each global surrogate loss model employing model-based sampling, we selected the learning rate that demonstrated the best performance from $\{0.01,0.05,0.1,0.5,1\}$. We run 10 experiments for each settings to ensure statistical significance. The experiments were performed on a Ryzen 7 5800X CPU and an RTX 3060 GPU with 64GB of RAM.
5.2 Results
Table 1 shows the normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ with standard error mean (SEM) for the inventory stock, budget allocation with 500 fake targets and portfolio optimization problems. For the surrogate loss models, we used 32 samples to learn the loss. We categorize the methods into four major training paradigms: two-stage, exact differentiation, local surrogate loss and global surrogate loss.
In the inventory stock problem, LODL-DQ, EGL-WMSE and EGL-DQ suffer from high regret, indicating that they do not provide informative gradients. LANCER and LCGLN perform well compared to others due to their expressiveness, as both use neural networks to learn the loss. For the budget allocation problem, we conduct four experiments with $\{0,5,50,500\}$ fake targets. Detailed results with different fake targets are available in Appendix C.1. For the hardest problem with 500 fake targets, the surrogate models tend to show better results than PFL or DFL methods. In the portfolio optimization problem, good prediction in parameter showed better decisions. This is evidenced by PFL and DFL showing better results compared to surrogate models. EGL-WMSE and LCGLN also demonstrated strong performance in our settings.
To reduce computational cost, it is crucial to train the surrogate loss with small sample sizes. We conducted experiments with varying sample sizes, as shown in the line plot in Figure 3. Note that the normalized test regret is better when lower. Since PFL, DFL and LANCER do not vary with changes in number of sample predictions, they show consistent results across all sample sizes. The surrogate models showed decreased regret as the sample size increased. At a sample size of 32, LCGLN outperformed the provided baselines. For detailed experimental results for each sample size and problem, please refer to Appendix C.2.
We also tested different numbers of fake targets, varying in $\{0,5,50,500\}$, to show how models perform on harder problems. We used 16 samples to train the loss model. Figure 4 presents the normalized test regret with SEM for global models across each experimental setting.
6 Conclusion
In this paper, we propose Locally Convex Global Loss Network (LCGLN), a global and general surrogate loss for DFL. LCGLN utilizes PCINN as a parametric surrogate to approximate task loss. We use PICNN to guarantee the surrogate to be convex near instances while maintaining a general non-convex structure globally. With LCGLN, users do not need to manually select the appropriate parametric function form for the task loss. Consequently, LCGLN can address general DFL problems with a single surrogate loss, regardless of the amount of observed data. We evaluated our method on three stochastic optimization problems, achieving better decision quality with fewer training samples compared to the state-of-the-art baselines. However, despite the expressive power of LCGLN, a limitation remains: achieving high decision quality requires careful selection of the samples. Thus, our future research will focus on identifying sampling strategies to improve the decision quality for surrogate loss models.
Appendix A Locally Convex Global Loss Network Details
A.1 Architecture Details
Here we show the architecture details of our Locally Convex Global Loss Network (LCGLN). The input is given as the pair of sampled prediction $\hat{y}$ and its corresponding instance $y$.
The bluish boxes represent the vectors that ensure convexity with respect to the input, while the reddish boxes correspond to the vectors that do not guarantee convexity. Yellow boxes denote the hidden layers, where the superscript inside the parentheses indicates the dimension, and the subscript identifies the layer number. For instance, $h_{i}^{(z)}$ is an $i$-th hidden layer with $h_{i}^{(z)}\in\mathbb{R}^{z}$.
We depict linear layers as trapezoids in gray when the dimensions change and as rectangles when they remain constant. The hatched box represents the non-negative linear layer. Details for operators are available in Figure 5.
The left side of the figure shows the architecture from input to the first hidden layer. The right side presents how hidden layers are composed. The upstream of the right side is a flow for reddish vector representations, while the downside is for bluish vectors. For the output layer, one may may simply let $z_{i+1}$ as a output vector and ignore $u_{i+1}$.
Figure 5: Architecture of Locally Convex Global Loss Network. Left figure represents the input to the first hidden layer. The right figure shows the hidden to hidden layers for non-convex vectors (reddish, upstream) and convex vectors (bluish, downstream). Output layer can simply be considered as $z_{i+1}$ without $u_{i+1}$.
Appendix B Experiment Details
B.1 Codes for Paper
The codes for experiments in our paper are available at the following link.
Codehttps://anonymous.4open.science/r/lcgln-5E21/README.md
B.2 Details in Inventory Stock Problem
The cost function is defined as:
$f(y,a)=c_{0}a+\frac{1}{2}q_{0}a^{2}+c_{b}[y-a]_{+}+\frac{1}{2}q_{b}([y-a]_{+})^{2}+c_{h}[a-y]_{+}+\frac{1}{2}q_{h}([a-y]_{+})^{2}$ (5)
where our objective is to minimize the expectation of cost function:
$\operatorname*{minimize}_{a}\quad\boldsymbol{E}_{y\sim p(y|x;\theta)}\left[f(y,a)\right].$ (6)
By the discrete distribution assumption of demand $y$, we can rewrite the optimization problem as:
$\operatorname*{minimize}_{a\in\mathbb{R},a_{b},a_{h}\in\mathbb{R}^{k}}$ $c_{0}a+\frac{1}{2}q_{0}a^{2}+\sum_{i=1}^{k}(p_{\theta})_{i}\left(c_{b}(a_{b})_{i}+\frac{1}{2}q_{b}(a_{b})_{i}^{2}+c_{h}(a_{h})_{i}+\frac{1}{2}q_{h}(a_{h})_{i}^{2}\right)$ (7)
subject to $d-a\boldsymbol{1}\leq a_{b},\;\,a\boldsymbol{1}-d\leq a_{h},\;\,a,a_{h},a_{b}\geq 0.$
where $(p_{\theta})_{i}\equiv p(y=d_{i}|x;\theta)$ for discrete demands $d_{1},\ldots,d_{k}$.
The parameter $c$ and $q$ each represent the coefficient of linear and quadratic costs, respectively. Specifically, $c_{0},q_{0}$ is related to the costs on the amount of product ordered, $c_{b},q_{b}$ and $c_{h},q_{h}$ is related to the amount of under and over ordered. We use the setting $(c_{0},q_{0},c_{b},q_{b},c_{h},q_{h})$ with $(10,2,30,14,10,2)$.
Appendix C Additional Results
C.1 Fake Target Variants Results for Budget Allocation
Here, we show the results of budget allocation with varying numbers of fake targets. We tested problems having $\{0,5,50,500\}$ fake targets, where the problem becomes more challenging as the fake targets increases. We use 16 samples for surrogate loss models. The results are presents in Table 2. The normalized test regret is lower the better, and we bold-lettered the results for each fake target experiments.
Table 2: The table presents normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ with standard error mean (SEM) tested on the budget allocation problem. The metric is lower the better with 0 meaning optimal. We tested for varying fake target numbers of $\{0,5,50,500\}$. Surrogate loss models were trained with 16 samples. The best-performing results for each number of fake targets are bold-lettered. We evaluate our LCGLN method against all baselines: PFL, DFL, and both local and global surrogate loss models.
Method # Fake Targets
0 5 50 500
PFL 0.128 $\pm$ 0.018 0.193 $\pm$ 0.028 0.242 $\pm$ 0.026 0.513 $\pm$ 0.016
DFL 0.064 $\pm$ 0.016 0.175 $\pm$ 0.026 0.377 $\pm$ 0.012 0.502 $\pm$ 0.011
LODL-DQ 0.133 $\pm$ 0.020 0.184 $\pm$ 0.017 0.326 $\pm$ 0.018 0.509 $\pm$ 0.008
LANCER 0.093 $\pm$ 0.013 0.185 $\pm$ 0.026 0.282 $\pm$ 0.031 0.490 $\pm$ 0.010
EGL-WMSE 0.189 $\pm$ 0.035 0.210 $\pm$ 0.033 0.340 $\pm$ 0.027 0.518 $\pm$ 0.007
EGL-DQ 0.068 $\pm$ 0.021 0.155 $\pm$ 0.037 0.289 $\pm$ 0.016 0.496 $\pm$ 0.016
LCGLN 0.090 $\pm$ 0.013 0.112 $\pm$ 0.024 0.272 $\pm$ 0.015 0.478 $\pm$ 0.007
C.2 Sample Size Variants Results
We present our experimental results using different sample sizes of $\{2,4,8,16,32\}$ for each problem and methodology. The results of normalized test regret for inventory stock, budget allocation, and portfolio optimization problem are in Table 3, 4, and 5 respectively. Note that the result of PFL and DFL remain unchanged regardless of the number of samples since they do not sample predictions. We bold-lettered the best performing results for each sample sizes.
Table 3: The table presents normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ with standard error mean (SEM) tested on the inventory stock problem. The metric is lower the better with 0 meaning optimal. We tested for varying sample sizes of $\{2,4,8,16,32\}$. The best-performing results for each sample size are bold-lettered. We evaluate our LCGLN method against all baselines: PFL, DFL, and both local and global surrogate loss models. LCGLN outperformed the baselines as the number of samples reached 32.
Method # Samples
2 4 8 16 32
PFL 0.242 $\pm$ 0.005 0.242 $\pm$ 0.005 0.242 $\pm$ 0.005 0.242 $\pm$ 0.005 0.242 $\pm$ 0.005
DFL 0.228 $\pm$ 0.002 0.228 $\pm$ 0.002 0.228 $\pm$ 0.002 0.228 $\pm$ 0.002 0.228 $\pm$ 0.002
LODL-DQ 0.371 $\pm$ 0.005 0.365 $\pm$ 0.003 0.373 $\pm$ 0.006 0.375 $\pm$ 0.005 0.378 $\pm$ 0.007
LANCER 0.182 $\pm$ 0.004 0.182 $\pm$ 0.004 0.182 $\pm$ 0.004 0.182 $\pm$ 0.004 0.182 $\pm$ 0.004
EGL-WMSE 0.383 $\pm$ 0.007 0.372 $\pm$ 0.004 0.371 $\pm$ 0.006 0.365 $\pm$ 0.003 0.371 $\pm$ 0.002
EGL-DQ 0.376 $\pm$ 0.005 0.368 $\pm$ 0.003 0.367 $\pm$ 0.002 0.366 $\pm$ 0.001 0.369 $\pm$ 0.007
LCGLN 0.237 $\pm$ 0.005 0.233 $\pm$ 0.004 0.224 $\pm$ 0.004 0.208 $\pm$ 0.007 0.174 $\pm$ 0.002
Table 4: The table presents normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ with standard error mean (SEM) tested on the budget allocation problem with 500 fake targets. The metric is lower the better with 0 meaning optimal. We tested for varying sample sizes of $\{2,4,8,16,32\}$. The best-performing results for each sample size are bold-lettered. We evaluate our LCGLN method against all baselines: PFL, DFL, and both local and global surrogate loss models. LCGLN outperformed the baselines as the number of samples reached 16.
Method # Samples
2 4 8 16 32
PFL 0.513 $\pm$ 0.016 0.513 $\pm$ 0.016 0.513 $\pm$ 0.016 0.513 $\pm$ 0.016 0.513 $\pm$ 0.016
DFL 0.502 $\pm$ 0.011 0.502 $\pm$ 0.011 0.502 $\pm$ 0.011 0.502 $\pm$ 0.011 0.502 $\pm$ 0.011
LODL-DQ 0.533 $\pm$ 0.008 0.523 $\pm$ 0.011 0.509 $\pm$ 0.011 0.509 $\pm$ 0.008 0.503 $\pm$ 0.020
LANCER 0.490 $\pm$ 0.010 0.490 $\pm$ 0.010 0.490 $\pm$ 0.010 0.490 $\pm$ 0.010 0.490 $\pm$ 0.010
EGL-WMSE 0.537 $\pm$ 0.008 0.529 $\pm$ 0.007 0.520 $\pm$ 0.006 0.518 $\pm$ 0.007 0.510 $\pm$ 0.013
EGL-DQ 0.520 $\pm$ 0.013 0.514 $\pm$ 0.015 0.499 $\pm$ 0.012 0.496 $\pm$ 0.016 0.492 $\pm$ 0.005
LCGLN 0.529 $\pm$ 0.016 0.496 $\pm$ 0.020 0.492 $\pm$ 0.008 0.478 $\pm$ 0.007 0.468 $\pm$ 0.009
Table 5: The table presents normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ with standard error mean (SEM) tested on the portfolio optimization problem. The metric is lower the better with 0 meaning optimal. We tested for varying sample sizes of $\{2,4,8,16,32\}$. The best-performing results for each sample size are bold-lettered. We evaluate our LCGLN method against all baselines: PFL, DFL, and both local and global surrogate loss models. LCGLN outperformed the baselines for all sample sizes.
Method # Samples
2 4 8 16 32
PFL 0.189 $\pm$ 0.002 0.189 $\pm$ 0.002 0.189 $\pm$ 0.002 0.189 $\pm$ 0.002 0.189 $\pm$ 0.002
DFL 0.187 $\pm$ 0.002 0.187 $\pm$ 0.002 0.187 $\pm$ 0.002 0.187 $\pm$ 0.002 0.187 $\pm$ 0.002
LODL-DQ 0.214 $\pm$ 0.003 0.212 $\pm$ 0.004 0.202 $\pm$ 0.002 0.195 $\pm$ 0.002 0.193 $\pm$ 0.002
LANCER 0.228 $\pm$ 0.006 0.228 $\pm$ 0.006 0.228 $\pm$ 0.006 0.228 $\pm$ 0.006 0.228 $\pm$ 0.006
EGL-WMSE 0.191 $\pm$ 0.002 0.190 $\pm$ 0.003 0.189 $\pm$ 0.002 0.188 $\pm$ 0.003 0.187 $\pm$ 0.001
EGL-DQ 0.260 $\pm$ 0.005 0.257 $\pm$ 0.004 0.257 $\pm$ 0.004 0.256 $\pm$ 0.003 0.256 $\pm$ 0.002
LCGLN 0.186 $\pm$ 0.000 0.185 $\pm$ 0.000 0.185 $\pm$ 0.000 0.185 $\pm$ 0.000 0.185 $\pm$ 0.000